OpenStudy (babybritt002):

Given that an = 3an - 1 + 1, where a1 = 2, find the 3rd-6th terms of the sequence.

OpenStudy (anonymous):

plug n=2 and find a2,then a3,a4,a5,a6

OpenStudy (anonymous):

i give a clue. a2=3 a1+1=3*2+1=7 similarly put n=3,4,5,6